# JEPA-Reasoner: Decoupling Latent Reasoning from Token Generation

### v1

@@ Abstract
While Joint-Embedding Predictive Architecture (JEPA) has emerged as a powerful architecture for learning rich latent representations, it fundamentally lacks generative abilities. Meanwhile, latent space reasoning attempts for Transformer models like COCONUT do improve performance, but they ultimately rely on token-by-token generation, which still accumulates compounding error and relies on context information to gain reasoning insights. To address these limitations, we propose JEPA-Reasoner, a novel JEPA model enhanced with generative ability that reasons in latent space. We augment it with a separate action-taker model, Talker, to produce human-readable sentences. Our approach demonstrates that decoupling latent space reasoning and token generation enables JEPA-Reasoner to produce mixed latent vectors that might lay the foundation for multi-threaded reasoning, while performing autoregressive generation with superior robustness to compounding error.
@@ 1 Introduction
The Joint-Embedding Predictive Architecture (JEPA) (i-jepa) has demonstrated strong performance in learning semantic world representations, exhibiting superior world understanding ability compared with traditional end-to-end generative models that work in pixel or token space. By predicting abstract representations in latent space, the JEPA architecture is able to filter irrelevant details and preserve essential information needed for prediction (i-jepa). Such architecture has been proven to be a viable approach for representation learning and foundation model development. Various JEPA implementations, including I-JEPA (i-jepa), V-JEPA2 (v-jepa2), and M3-JEPA (m3-jepa), have shown success across various modalities and downstream tasks.
@@
However, JEPA models are inherently non-generative (m3-jepa) because of their objective: filling missing information in the current state (i-jepa), rather than generating new content. Besides, the predictor often requires a detailed, predetermined target state or task instruction (e.g., V-JEPA 2-AC (v-jepa2)), which is often unavailable in tasks demanding long-term planning and step-by-step reasoning. This feature limits the application of the broad knowledge in JEPA models to generative tasks.
@@
Furthermore, while traditional token-level autoregressive models have sequential reasoning capabilities, their token-by-token generation process is prone to compounding errors. Even if the predicted probabilities of tokens are weighted and combined, LLMs still cannot go beyond single-threaded reasoners (llms-reason-singel-thread). Although several previous research have explored latent space reasoning for Transformer models, the end-to-end generation goal of a single coupled model limited the full potential of latent space reasoning, while also making training complex and inefficient (coconut).
@@
To address these limitations, we propose JEPA-Reasoner, a novel decoupled architecture that utilizes separate models for reasoning and token generation. The reasoning model, JEPA-Reasoner, transforms the JEPA framework from a target-conditioned system into an autoregressive generative model. Operating entirely within the continuous, normalized latent space, JEPA-Reasoner focuses solely on latent space reasoning, offloading the token generation task to its action-taker module: Talker. In contrast to existing latent space reasoning solutions like COCONUT, the division of objective frees the reasoner from expression burden and enables continuous latent guidance (compared with non-continuous latent guidance that relies on context information to retrieve latent reasoning results (coconut)) during the token generation process.
@@
Our key insight is that performing pure reasoning in an abstract, continuous, and normalized latent space without token generation burden allows for the construction of high-level reasoning chains that carry rich semantic information throughout the autoregressive process while correctly ignoring irrelevant details or distracting information. Our empirical experiments show further improved abilities beyond that: by operating in the latent space, JEPA-Reasoner can maintain multiple hypotheses during the reasoning process simultaneously and mitigate the catastrophic error propagation associated with discrete token sampling.
@@ 2 Related Work
JEPA (i-jepa) introduced a framework that makes predictions in representation space. It utilizes self-supervised training to learn latent states that are not directly human-readable. A predictor module was trained to predict the target state based on encoded inputs. Multiple variants of this architecture have extended the JEPA family to various modalities and downstream tasks (i-jepa). However, these models are non-generative. Attempts to make JEPA generative, such as D-JEPA (d-jepa), utilize the learned representations to condition diffusion models for data generation (e.g., text to images, text to audio) but still failed to enable sequential reasoning or planning within the JEPA framework itself. In contrast, our key innovation is to adapt the core JEPA objective for autoregressive latent-space generation.
@@
Previous work on latent space reasoning mainly focuses on looping hidden states, either through horizontal autoregression like COCONUT (coconut), or vertical recurrent depth scaling (recurrent-latent-reasoning). However, these paradigms utilize a single coupled model for both latent reasoning and token generation, ignoring the mismatch between the two tasks: one requires high-level global planning, decision-making, and choice tracing, while the other requires local correctness in grammar and fluency. Besides, whether iterating over tokens or deepening the computation per token, both paradigms remain bound by causal masking, meaning early token generation cannot be guided by future reasoning states. Considering these limitations, our key innovation is to decouple latent space reasoning and token generation. By generating a complete latent chain first, we enable consistent latent guidance where all tokens are generated from the full reasoning trajectory. This produces an answer with higher quality that is less prone to error propagation. Additionally, the decoupled design enables efficient optimization with a single forward pass in latent space, unlike coupled models like COCONUT or Recurrent Depth transformers, which require multiple synchronized passes or complex recurrent unrolling.
@@
Modern Transformer models conduct token-level prediction in an autoregressive manner. While proven powerful on various tasks, this approach is known to suffer from compounding errors in long-horizon tasks. Techniques like Chain-of-Thought (cot) improve reasoning by generating intermediate steps, but still operate at the token level. JEPA-Reasoner aims to improve robustness of autoregressive generation by moving the reasoning process into a continuous, abstract latent space, reducing the impact of localized errors.
@@ 3.1 JEPA-Reasoner
JEPA-Reasoner follows the JEPA philosophy, containing an embedding layer as a textual token encoder and modified Transformer blocks for the predictor, since the Transformer has proven its strong ability in sequence modeling. After the modified Transformer blocks, we applied a hybrid normalization layer (RMS and L2 normalization). We utilize L2 normalization to prevent exploding magnitude caused by residual connections. In the modified Transformer block, we apply a non-learnable QK-Norm (qk-norm) to make it more numerically stable.
@@ 3.2 Action-Taker Model
In this experiment, there is only one action-taker model: Talker. The Talker model is a standard Transformer-based model trained independently. We designed two variants of Talker: Mono-Talker and Dual-Talker. Detailed information about the components of the two Talkers is shown in Table 1. Mono-Talker does not have an embedding layer or encoders, it only has decoders. Mono-Talker is designed for reconstruction tasks that do not require context information, receiving latent vectors from JEPA-Reasoner and constructs the complete token sequence in one forward pass. Dual-Talker is designed for context-aware reconstruction, usually necessary in natural language tasks. It has an embedding layer, encoders, and decoders. The embedding layer is used for encoding previously determined outputs of JEPA-Reasoner that contain contextual information, while the encoder blocks receive latent vectors from JEPA-Reasoner as input. The decoders generate tokens autoregressively conditioned on previous tokens, with continuous latent guidance from the output of encoders. However, Dual-Talker was trained for reconstruction rather than generation, as our ablation study (Appendix C) showed that Talker is critically dependent on the Reasoner’s output. During training, the JEPA-Reasoner is frozen. Talker receives the sequence of latent vectors and is trained to reconstruct the corresponding token sequence using standard cross-entropy loss.
@@ 4 Training Procedure
The training process of JEPA-Reasoner consists of two main phases. The first stage is pretraining that teaches basic knowledge (e.g., grammar and commonsense) to the model. The second stage is self-supervised training (SST), which adapts the model to perform consistent latent space reasoning.
@@ 4.1 Pretraining
The model is trained as a standard decoder-only Transformer on the next-token prediction task in a teacher-forcing way. We employ tied word embeddings that share the weight of the embedding layer with a temporary LM head. The LM head is only used in pretraining and is removed after pretraining is finished. The L2 normalization layer is disabled in the pretraining phase to enable simple and direct reuse of current Transformer training recipes. Considering that tied word embedding encourages $\mathbf{W}_{Embed}\cdot\mathbf{W}_{Embed}^{T}=I$ and $\mathbf{v}_{pred}\cdot\mathbf{v}_{embed}=\|\mathbf{v}_{pred}\|\cdot\|\mathbf{v}_{embed}\|\cdot\cos(\theta)$, the tied-embedding approach will indirectly encourage angular alignment between predicted vectors and embedding vectors, which facilitates the subsequent transition from token-level to latent-level prediction.
@@ 4.2 Self-Supervised Training
The SST phase adapts the pretrained model for making consistent predictions in the continuous latent space. Since the model is fully transforming into a latent space reasoner instead of a token generator, the ability to produce correct logits no longer matters. Considering this, we apply similar self-supervised training as Meta’s JEPA series (i-jepa) in this stage. Without the need for autoregressively generating final token outputs to compute a loss, self-supervised training enables efficient parallel training compared with COCONUT (coconut).
Objective and Methodology.
The temporary LM head in the pretraining stage was discarded and L2 normalization layer was restored. The model is now optimized to predict the latent representation of the next sequence segment, with a consistent dimensional semantics as what the embedding layer produces. We switch to scaled cosine distance loss, aligning with the L2 normalization strategy used to ensure stability during autoregressive looping and focusing the learning on angular differences:
$\mathcal{L}(\theta,\theta^{\prime})=k-k\cdot\text{cos}(h_{\text{pred}}(\theta),h_{\text{target}}(\theta^{\prime}))$ (1)
where $k$ is the scalar, $h_{\text{pred}}$ is the predicted latent vector by the Reasoner (parameters $\theta$), and $h_{\text{target}}$ is the target latent vector from the EMA encoder (parameters $\theta^{\prime}$). In our empirical tests, we find that normal cosine distance loss failed to support enough optimization when the loss is small. We tested a series of $k$ values, and chose $k=4$ in our experiments (Refer to Appendix D for more detail).
Target Generation.
The weight updating method of the target embedding layer is different from that of the input embedding layer. The input embedding layer always applies the latest weights, while the target embedding layer utilizes exponential moving average to generate its weights from the historical weights of the input embedding layer. We applied a high momentum value of 0.98 to prevent rank collapse in the embedding layer while ensuring enough space to adjust for angular alignment.
@@ 5.2 Robustness to Error Perturbation
In the following sections, we demonstrate that decoupling reasoning chain generation from token production enables superior robustness under noisy conditions. While coupled models must simultaneously maintain reasoning coherence and produce correct tokens, our decoupled approach allows the reasoning model to focus solely on maintaining logical consistency in latent space, exhibiting better generation quality. This section contains two experiments that focus on two different sources of errors: token level error and latent space noise.
@@ 5.2.1 Experiment Methods
Robustness Test for Token Level Error
@@
Robustness Test for Latent Space Error
To evaluate the robustness of the JEPA-Reasoner model on perturbations in latent space, we compare the performance of JEPA-Reasoner and the coupled continuous reasoning model COCONUT on multi-step completion tasks. In this experiment, we let the COCONUT model autoregressively generate 4 latent vectors first, followed by 4 tokens. For JEPA-Reasoner, we simply let it generate 8 latent vectors and use the Talker module to reconstruct 8 tokens. For both models, we add Gaussian noise to the generated latent vector at each step with $\mu=0$ and $\sigma$ ranging from 0% to 15% of the maximum value in the model’s output. Accuracy is calculated across the last 4 tokens with the exact match metric.
@@ 5.2.3 Model Configurations and Training Methods
We first pretrain the Transformer models on CFG data using cross-entropy loss in token space. Since the pretraining methods of Transformer, COCONUT, and JEPA-Reasoner are identical, subsequent trainings are based on the same pretraining checkpoint [^2] Due to architectural differences, we initialize JEPA-Reasoner models using only the first $N$ blocks from the pretrained Transformer, where $N$ matches the JEPA-Reasoner’s block count.. We conduct post-training to obtain the Transformer baseline $T$. COCONUT models are trained to first predict 4 hidden states, then generate 4 tokens. Cross-entropy loss is computed between the output logits and the target sequence, excluding the hidden state positions. We follow the training method mentioned in Section 4 to train the JEPA-Reasoner and Dual Talker model $R$. For all training processes mentioned above, we apply identical hyperparameters (with learning rate of $1\times 10^{-4}$, effective batch size of 128, and context length of 1024) and train until their loss stabilizes, then checkpoints of best performance were chosen as the representative.
@@ B.1 CFG Rules and Sample
CFG can hierarchically produce highly structured expressions by replacing non-terminal symbols at each level with next-level symbols following a production rule, as shown in Figure 4. A sequence of terminal symbols is considered to be valid if it can be transformed back to the root symbol with dynamic programming and the given production rule. The recursive structure and local ambiguity of CFG sequences enable them to model the rich and recursive structure in languages, including grammar and logic. We designed our own CFG following the method used by (allen-physics). The production rule used in our experiments is a five-level CFG production rule set featuring three terminal symbols with 3 or 4 rule lengths at each level, which typically generates long (typically 600 to 700 symbols per sample) and locally ambiguous sequences. A visualization of the production rule used in this experiment can be seen in Figure 5.
@@
Figure 5: CFG production rule used to generate training and test samples in Section 5.2. This rule gives sequence lengths ranging from about 600 symbols to 700 symbols.

### v2

@@ Abstract
While Joint-Embedding Predictive Architecture (JEPA) has emerged as a powerful architecture for learning rich latent representations, it fundamentally lacks generative abilities. Meanwhile, current latent reasoning models remain limited by the token-by-token generation paradigm, which suffers from compounding errors and heavy context dependency. To address these limitations, we proposed JEPA-Reasoner, a novel JEPA-based architecture enhanced with generative ability for latent reasoning. We augment this architecture with a separate action-talker model, Talker, to reconstruct human-readable text from latent representations produced by the JEPA-Reasoner. Our work demonstrated that decoupling latent-space reasoning from token production enables JEPA-Reasoner to produce mixed latent vectors, laying a foundation for multi-threaded reasoning and achieving superior robustness against compounding errors in autoregressive generation.
@@ 1 Introduction
The Joint-Embedding Predictive Architecture (JEPA) (i-jepa) has demonstrated strong performance in learning semantic world representations, exhibiting a superior ability to understand the world compared with traditional end-to-end generative models that operate in pixel- or token-space. By predicting abstract representations in latent space, the JEPA architecture can filter irrelevant details and preserve essential information needed for prediction (i-jepa). Such architecture has been proven to be a viable approach for representation learning and foundation model development. Various JEPA implementations, including I-JEPA (i-jepa), V-JEPA2 (v-jepa2), and M3-JEPA (m3-jepa), have shown success across various modalities and downstream tasks.
@@
However, JEPA models are inherently non-generative (m3-jepa) because of their objective: filling missing information in the current state (i-jepa), rather than generating new content. Furthermore, the predictor often requires a detailed, predetermined target state or task instruction (e.g., V-JEPA 2-AC (v-jepa2)), which is typically unavailable in tasks that demand long-term planning and step-by-step reasoning. This feature limits the application of the broad knowledge in JEPA models to generative tasks.
@@
Furthermore, while traditional token-level autoregressive models have sequential reasoning capabilities, their token-by-token generation process is prone to compounding errors. Even if the predicted probabilities of tokens are weighted and combined, LLMs still cannot go beyond single-threaded reasoners (llms-reason-singel-thread). Although several previous studies have explored latent space reasoning for Transformer models, the end-to-end generation goal of a single coupled model has limited the full potential of latent space reasoning and made training complex and inefficient (coconut).
@@
To address these limitations, we propose JEPA-Reasoner, a novel decoupled architecture that utilizes separate models for reasoning and token generation. The reasoning model, JEPA-Reasoner, transforms the JEPA framework from a target-conditioned system into an autoregressive generative model. Operating entirely within the continuous, normalized latent space, JEPA-Reasoner focuses solely on latent-space reasoning, offloading the token-generation task to its action-taker module, Talker. In contrast to existing latent space reasoning solutions like COCONUT, the division of objective frees the reasoner from expression burden and enables continuous latent guidance (compared with non-continuous latent guidance that relies on context information to retrieve latent reasoning results (coconut)) during the token generation process.
@@
Our key insight is that performing pure reasoning in an abstract, continuous, and normalized latent space without the burden of token generation enables the construction of high-level reasoning chains that carry rich semantic information throughout the autoregressive process while correctly ignoring irrelevant details or distracting information. Our empirical experiments demonstrate further improvements beyond that: by operating in the latent space, JEPA-Reasoner can maintain multiple hypotheses simultaneously during the reasoning process and mitigate catastrophic error propagation associated with discrete token sampling.
@@ 2 Related Work
JEPA (i-jepa) introduced a framework that makes predictions in representation space. It utilizes self-supervised training to learn latent states that are not directly human-readable. A predictor module was trained to predict the target state based on encoded inputs. Multiple variants of this architecture have extended the JEPA family to various modalities and downstream tasks (i-jepa). However, these models are non-generative. Attempts to make JEPA generative, such as D-JEPA (d-jepa), use learned representations to condition diffusion models for data generation (e.g., text-to-image, text-to-audio), but they still fail to enable sequential reasoning or planning within the JEPA framework itself. In contrast, our key innovation is to adapt the core JEPA objective for autoregressive latent-space generation.
@@
Previous work on latent space reasoning mainly focuses on looping hidden states, either through horizontal autoregression like COCONUT (coconut), or vertical recurrent depth scaling (recurrent-latent-reasoning). However, these paradigms utilize a single coupled model for both latent reasoning and token generation, ignoring the mismatch between the two tasks: one requires high-level global planning, decision-making, and choice tracing, while the other requires local correctness in grammar and fluency. Besides, whether iterating over tokens or deepening the computation per token, both paradigms remain bound by causal masking, meaning early token generation cannot be guided by future reasoning states. Given these limitations, our key innovation is to decouple latent-space reasoning and token generation. By generating a complete latent chain first, we enable consistent latent guidance, with all tokens generated from the whole reasoning trajectory. This produces a higher-quality answer that is less prone to error propagation. Additionally, the decoupled design enables efficient optimization with a single forward pass in latent space, unlike coupled models like COCONUT or Recurrent Depth transformers, which require multiple synchronized passes or complex recurrent unrolling.
@@
Modern Transformer models conduct token-level prediction in an autoregressive manner. While proven powerful on various tasks, this approach is known to suffer from compounding errors in long-horizon tasks. Techniques like Chain-of-Thought (cot) improve reasoning by generating intermediate steps, but still operate at the token level. JEPA-Reasoner aims to enhance robustness of autoregressive generation by moving the reasoning process into a continuous, abstract latent space, reducing the impact of localized errors.
@@ 3.1 JEPA-Reasoner
JEPA-Reasoner follows the JEPA philosophy, containing an embedding layer as a textual token encoder and modified Transformer blocks for the predictor, since the Transformer has proven its strong ability in sequence modeling. After the modified Transformer blocks, we applied a hybrid normalization layer (RMS and L2 normalization). We utilize L2 normalization to prevent exploding magnitude caused by residual connections. In the modified Transformer block, we apply a non-learnable QK-Norm (qk-norm) to ensure numerical stability.
@@ 3.2 Action-Taker Model
In this experiment, there is only one action-taker model: Talker. The Talker model is a standard Transformer-based model trained independently. We designed two variants of Talker: Mono-Talker and Dual-Talker. Detailed information about the components of the two Talkers is shown in Table 1. Mono-Talker consists of decoders only, without an embedding layer or encoders. Mono-Talker is designed for reconstruction tasks that do not require context information, receiving latent vectors from JEPA-Reasoner and constructs the complete token sequence in one forward pass. Dual-Talker is designed for context-aware reconstruction, usually necessary in natural language tasks. It has an embedding layer, encoders, and decoders. The embedding layer is used for encoding previously determined outputs of JEPA-Reasoner that contain contextual information, while the encoder blocks receive latent vectors from JEPA-Reasoner as input. The decoders generate tokens autoregressively conditioned on previous tokens, with continuous latent guidance from the output of encoders. However, Dual-Talker was trained for reconstruction rather than generation, as our ablation study (Appendix C) showed that Talker is critically dependent on the Reasoner’s output. During training, the JEPA-Reasoner is frozen. Talker receives the sequence of latent vectors and is trained to reconstruct the corresponding token sequence using standard cross-entropy loss.
@@ 4 Training Procedure
The training process of JEPA-Reasoner consists of two main phases. (1) Pretraining, to acquire linguistic competence and world knowledge, and (2) Self-Supervised Training (SST), to adapt the model for consistent reasoning in the continuous latent space.
@@ 4.1 Pretraining
The model is trained as a standard decoder-only Transformer on the next-token prediction task via teacher forcing. To facilitate the transition to the subsequent SST phase, we introduce specific architectural changes:
  • Tied Embeddings: We share weights between the embedding layer and a temporary Language Model (LM) head. This LM head is discarded after pretraining.
  • L2 Normalization: We disable the L2 normalization layer during this phase to maintain compatibility with standard Transformer training recipes.
The use of tied embeddings also indirectly facilitates angular alignment between the predicted vectors and the embedding vectors since it encourages $\mathbf{W}_{Embed}\cdot\mathbf{W}_{Embed}^{T}=I$ and $\mathbf{v}_{pred}\cdot\mathbf{v}_{embed}=\|\mathbf{v}_{pred}\|\cdot\|\mathbf{v}_{embed}\|\cdot\cos(\theta)$. This implicit angular alignment creates a smoother transition from token-level prediction to latent-level prediction in the subsequent SST phase.
@@ 4.2 Self-Supervised Training
The SST phase transforms the model from a token generator into a pure latent-space reasoner. We adopt a self-supervised approach inspired by the JEPA series (i-jepa). Unlike methods such as COCONUT (coconut), our approach avoids the computational overhead of autoregressively generating final tokens or looping latent vectors multiple times for loss computation, enabling more efficient parallel training.
Optimization Objective.
The model is now optimized to predict the latent representation of the next sequence segment. We discard the temporary LM head and restore the L2 normalization layer. We apply a similar self-supervised training strategy proposed in Meta’s JEPA series (i-jepa) and switch to scaled cosine distance loss, aligning with the L2 normalization strategy to ensure stability during autoregressive looping and encouraging the model to learn angular differences:
$\mathcal{L}(\theta,\theta^{\prime})=k\cdot\left(1-\cos(h_{\text{pred}}(\theta),h_{\text{target}}(\theta^{\prime}))\right)$ (1)
where $k$ is a scalar, $h_{\text{pred}}$ is the latent vector predicted by the Reasoner (parameters $\theta$) and $h_{\text{target}}$ is the target vector from the EMA encoder (parameters $\theta^{\prime}$). We adopt a scaling factor of $k=4$ as we empirically observed that standard cosine distance yields insufficient gradients when the loss is small. (See Appendix D for more discussions about scaled cosine distance loss).
Target Generation (EMA).
To prevent rank collapse while allowing angular adjustment, we employ an asymmetric weight update strategy:
  1. 1.
    Input Embedding (Online): Updated via standard backpropagation.
  2. 2.
    Target Embedding (Target): Updated via exponential moving average (EMA) of the input embedding weights. We applied a high momentum value of 0.98 to prevent rank collapse in the embedding layer while ensuring enough space to adjust for angular alignment.
@@ 5.2 Robustness to Error Perturbation
We demonstrate that decoupling reasoning generation from token production yields superior robustness. While coupled models must simultaneously maintain reasoning coherence and produce correct tokens, our decoupled approach allows the reasoning model to focus solely on logical consistency in the latent space, demonstrating superior generation quality and robustness to noisy input. We validate this advantage through two experiments targeting token-level errors and latent space noise, respectively.
@@ 5.2.1 Experiment Methods
Robustness Test for Token Level Error.
@@
Robustness Test for Latent Space Error.
To evaluate the robustness of the JEPA-Reasoner model on perturbations in latent space, we compare the performance of JEPA-Reasoner and the coupled continuous reasoning model COCONUT on multi-step completion tasks. In this experiment, COCONUT generates four latent vectors followed by four tokens, whereas JEPA-Reasoner generates eight latent vectors and reconstructs eight tokens via the Talker module. We inject Gaussian noise at each step ($\mu=0$; $\sigma\in[0\%,15\%]$ relative to the maximum output value) and measure exact match accuracy on the final four tokens.
@@ 5.2.3 Model Configurations and Training Methods
We first pretrain Transformer models on CFG data with cross-entropy loss, then pick the checkpoint with optimal accuracy on CFG completion task. This unified checkpoint serves as the initialization for all subsequent models. [^2] The JEPA-Reasoner is initialized using the first $N$ blocks of the pretrained Transformer to match its layer count. We then post-train the Transformer to obtain baseline $T$. COCONUT models are trained to predict a sequence of four hidden states followed by four tokens, with loss computed exclusively on the target tokens. The JEPA-Reasoner and Dual Talker model $R$ are trained following the methodology in Section 4. All training phases use identical hyperparameters (learning rate $1\times 10^{-4}$, effective batch size 128, context length 1024) and all models are trained until their loss stabilizes. The best-performing checkpoint is selected as the representative for each architecture.
@@ B.1 CFG Rules and Sample
CFG can hierarchically produce highly structured expressions by replacing non-terminal symbols at each level with next-level symbols following a production rule, as shown in Figure 4. A sequence of terminal symbols is considered to be valid if it can be transformed back to the root symbol with dynamic programming and the given production rule. The recursive structure and local ambiguity of CFG sequences enable them to model the rich and recursive structure in languages, including grammar and logic. We designed our own CFG following the method used by (allen-physics). The production rule used in our experiments is a five-level CFG production rule set featuring three terminal symbols with 3 or 4 rule lengths at each level, which typically generates long (typically 600 to 700 symbols per sample) and locally ambiguous sequences.